OpenStudy (anonymous):

[6.02] and [6.03] What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2

OpenStudy (anonymous):

use substitution to solve each problem

OpenStudy (anonymous):

so for the first one, 4x+8y=8

OpenStudy (anonymous):

4x=8-8y

OpenStudy (anonymous):

x=2-2y

OpenStudy (anonymous):

and the 4(2-2y)+8y=8

OpenStudy (anonymous):

so 8-8y+8=8

OpenStudy (anonymous):

8-8y=0

OpenStudy (anonymous):

-8y=-8

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

test it in both to see if it works, and then solve for x

OpenStudy (anonymous):

ohk thank you